# Towards Transferable Defense Against Malicious Image Edits

### v1

@@ I Introduction
Figure 1: Qualitative comparison demonstrating TDAE’s enhanced cross-model transferability. Both original images are immunized on StableDiffusion v3
to resist the edit prompt “Add a lion”. The top row utilizes the SA [34] method, while the bottom row employs SA+TDAE. Although the SA method shows limited effectiveness when transferred to the unseen InstructPix2Pix model, SA+TDAE successfully disrupts the malicious edit on both the source or surrogate model, StableDiffusion v3, and the unseen model, InstructPix2Pix, showcasing superior generalization.
In this paper, we propose Transferable Defense Against Malicious Image Edits (TDAE), a novel bimodal framework that enhances adversarial robustness by synergistically optimizing both image and text modalities.. Unlike traditional methods, TDAE specifically enhances the transferability of immunized images, which consists of two key parts: i.e., FlatGrad Defense Mechanism (FDM) and Dynamic Prompt Defense (DPD). Inspired by [37], we propose FDM by minimizing the expectation of gradient norms across perturbed images at the image level for enhancing immunized image transferability. FDM replaces the conventional expectation minimization with direct optimization of the maximum gradient norm under local perturbations, eliminating sampling requirements while maintaining robust defense effectiveness against unseen models. This reformulation achieves substantial efficiency gains without sacrificing effectiveness. Complementing FDM, we introduce DPD at the text level, leveraging insights from multi-prompt optimization [38]. DPD employs an iterative process to refine text embeddings, ensuring that edits on immunized images align with those of the original images. Concurrently, it updates immunized images using these optimized text embeddings, creating a dual-loop optimization that generates immunized images by seeking a broader set of immunity-enhancing features. Broader features bring potential transferability to unseen models, leading to better cross-model immunity effectiveness. It should be noted that our TDAE is a plug-and-play module which can be integrated into any image immunization method for improving cross-model transferability. In summary, we make the following contributions:
@@ II Related Work
This section reviews literature related to our proposed TDAE, covering three primary areas. These include advancements in diffusion-model-based image editing, existing image immunization techniques, and methodologies for enhancing the transferability of adversarial attacks.
@@ II-B Image Immunization
The rapid evolution of diffusion-based generative models has significantly lowered the barrier to malicious image manipulation, thereby accentuating the urgent need for robust countermeasures. Existing image immunization techniques address this challenge by strategically embedding targeted, often imperceptible, adversarial perturbations into original images [45, 29, 28, 46, 47, 48] to disrupt unauthorized modifications. For diffusion-based image editing models, Salman et al. [32] introduce two pioneering approaches named PGE and PGD. The PGE [32] method perturbs the VAE’s latent distribution, offering simplicity but limited protection, while PGD [32] disrupts the entire forward denoising process for stronger defense at a high computational cost. Building on these initial efforts, subsequent research explore more nuanced strategies. To mitigate high computational overheads, Lo et al. [34] propose selectively targeting text-salient regions via cross-attention layers. Concurrently, Mist et al. [35] incorporates textural losses to prevent unauthorized style mimicry by minimizing the distance in the latent space between the protected image and a pre-defined target image. Further efforts aim at enhancing attack effectiveness and efficiency. For example, Xue et al. [33] leverages score distillation sampling, a plug-and-play strategy that improves efficiency by approximating the gradient calculation. Zheng et al. [36] investigates targeted attacks, which works by minimizing the distance between the editing model’s output and a pre-defined target. Despite significant progress, a key challenge persists: images immunized against specific models often lose effectiveness against unseen or architecturally different models. This underscores the critical need for more generalized, model-agnostic immunization strategies.
@@ IV-A Experimental Setup
For intra-model evaluations, immunized images are generated by applying adversarial perturbations crafted against a specific source editing model. The effectiveness of these immunized images is then evaluated by attempting to edit them using the same source model. These intra-model assessments are conducted on INS, SD3, and SD14 respectively. This setting evaluates the defense capability of an immunization method against the model it is directly optimized for.
@@ IV-B Main Results
Qualitative comparisons in Fig. 4 further substantiate TDAE’s superior cross-model transferability. Baseline methods often yield edited images resembling benign edits of originals, while TDAE-enhanced methods consistently induce severe editing failures with substantial semantic distortions. For instance, during the INS to SD3 transfer with the “use more vibrant colors” prompt, TDAE-enhanced images significantly deviate from expectations or introduce unintended background alterations despite partial edit success. The PGE+FDM method notably degrades the image into abstract color fields. A similar pattern is observed in the SD3 to INS transfer for “turn the fisheye lens into a pinhole”. While the immunized images generated by the baseline method exhibit slight distortions during editing, they generally achieve the desired "turn the fisheye lens into a pinhole" effect. In contrast, although the TDAE-enhanced method also successfully implements the "turn the fisheye lens into a pinhole" edit, the overall color palette of the image dramatically shifts from blue to a dark gray. This consistent induction of severe distortions, semantic incongruities or undesirable global changes demonstrates TDAE’s robust capability to disrupt malicious edits on unseen models.

### v2

# arXiv:2512.14341v2[cs.CV] 02 Mar 2026

@@ I Introduction
Figure 1: Qualitative comparison demonstrating TDAE’s enhanced cross-model transferability. Both original images are immunized on StableDiffusion v3 to resist the edit prompt “Add a lion”. The top row utilizes the SA [34] method, while the bottom row employs SA+TDAE. Although the SA method shows limited effectiveness when transferred to the unseen InstructPix2Pix model, SA+TDAE successfully disrupts the malicious edit on both the source or surrogate model, StableDiffusion v3, and the unseen model, InstructPix2Pix, showcasing superior generalization.
In this paper, we propose Transferable Defense Against Malicious Image Edits (TDAE), a novel bimodal framework that enhances adversarial robustness by synergistically optimizing both image and text modalities. Unlike traditional methods, TDAE specifically enhances the transferability of immunized images, which consists of two key parts: FlatGrad Defense Mechanism (FDM) and Dynamic Prompt Defense (DPD). Inspired by [37], we propose FDM by minimizing the expectation of gradient norms across perturbed images at the image level for enhancing immunized image transferability. FDM replaces the conventional expectation minimization with direct optimization of the maximum gradient norm under local perturbations, eliminating sampling requirements while maintaining robust defense effectiveness against unseen models. This reformulation achieves substantial efficiency gains without sacrificing effectiveness. Complementing FDM, we introduce DPD at the text level, leveraging insights from multi-prompt optimization [38]. DPD employs an iterative process to refine text embeddings, ensuring that edits on immunized images align with those of the original images. Concurrently, it updates immunized images using these optimized text embeddings, creating a dual-loop optimization that generates immunized images by seeking a broader set of immunity-enhancing features. Broader features bring potential transferability to unseen models, leading to better cross-model immunity effectiveness. It should be noted that our TDAE is a plug-and-play module which can be integrated into any image immunization method for improving cross-model transferability. In summary, we make the following contributions:
@@ II Related Work
This section reviews literature related to our proposed TDAE, covering three primary areas. These include advancements in diffusion-model-based image editing, image immunization techniques, and methodologies for enhancing the transferability of adversarial attacks.
@@ II-B Image Immunization
The rapid evolution of diffusion-based generative models has significantly lowered the barrier to malicious image manipulation, thereby accentuating the urgent need for robust countermeasures. Existing image immunization techniques address this challenge by strategically embedding targeted, often imperceptible, adversarial perturbations into original images [45, 29, 28, 46, 47, 48] to disrupt unauthorized modifications. For diffusion-based image editing models, Salman et al. [32] introduce two pioneering approaches named PGE and PGD. The PGE [32] method perturbs the VAE’s latent distribution, offering simplicity but limited protection, while PGD [32] disrupts the entire forward denoising process for stronger defense at a high computational cost. Building on these initial efforts, subsequent research explores more nuanced strategies. To mitigate high computational overheads, Lo et al. [34] propose selectively targeting text-salient regions via cross-attention layers. Concurrently, Mist et al. [35] incorporate textural losses to prevent unauthorized style mimicry by minimizing the distance in the latent space between the protected image and a pre-defined target image. Further efforts aim at enhancing attack effectiveness and efficiency. For example, Xue et al. [33] leverage score distillation sampling, a plug-and-play strategy that improves efficiency by approximating the gradient calculation. Zheng et al. [36] investigate targeted attack, which works by minimizing the distance between the editing model’s output and a pre-defined target. Despite significant progress, a key challenge persists: images immunized against specific models often lose effectiveness against unseen or architecturally different models. This underscores the critical need for more generalized, model-agnostic immunization strategies.
@@ IV-A Experimental Setup
For intra-model evaluations, immunized images are generated by applying adversarial perturbations crafted against a specific source editing model. The effectiveness of these immunized images is then evaluated by attempting to edit them using the same source model. These intra-model assessments are conducted on INS, SD3, and SD14, respectively. This setting evaluates the defense capability of an immunization method against the model it is directly optimized for.
@@ IV-B Main Results
Qualitative comparisons in Fig. 4 further substantiate TDAE’s superior cross-model transferability. Baseline methods often yield edited images resembling benign edits of originals, while TDAE-enhanced methods consistently induce severe editing failures with substantial semantic distortions. For instance, during the INS to SD3 transfer with the “use more vibrant colors” prompt, TDAE-enhanced images significantly deviate from expectations or introduce unintended background alterations despite partial edit success. The PGE+FDM method notably degrades the image into abstract color fields. A similar pattern is observed in the SD3 to INS transfer for “turn the fisheye lens into a pinhole”. While the immunized images generated by the baseline method exhibit slight distortions during editing, they generally achieve the desired “turn the fisheye lens into a pinhole” effect. In contrast, although the TDAE-enhanced method also successfully implements the “turn the fisheye lens into a pinhole” edit, the overall color palette of the image dramatically shifts from blue to a dark gray. This consistent induction of severe distortions, semantic incongruities or undesirable global changes demonstrates TDAE’s robust capability to disrupt malicious edits on unseen models.